OpenStudy (beef123):

The value of -3xy, for x = 2 and y = 3, is -18. True or False

OpenStudy (seratul):

Just plug it into the equation -3 * 2 = -6 -6 * 3 = -18 So it is true.

OpenStudy (beef123):

Thanks alot